OpenStudy (thedj4jc):

NEED HELP QUICK!!!!!!! Given: X = r + 2, Y = 2r - 9, and Z = r2 + 17r + 30. Simplify [X · Y - Z] ÷ X. -r - 24 r - 24 2r2 - 6r - 33 2r2 + 6r + 33

OpenStudy (thedj4jc):

[R + 2 * 2R - 9 - R^2 + 17R + 30] / R + 2 That's how I'd write it out, right?

OpenStudy (thedj4jc):

@iGreen

OpenStudy (thedj4jc):

@robtobey

OpenStudy (thedj4jc):

@icecreamman

OpenStudy (anonymous):

Ooh this looks fun

OpenStudy (anonymous):

\[\frac{ (x \times y - z) }{ x }\] this is your problem right (just making sure)

OpenStudy (thedj4jc):

I'm not sure. I'm supposed to simplify the given equation, [X * Y - Z] / X

OpenStudy (thedj4jc):

Yes, pretty much. Sorry, had to write it out to notice it.

OpenStudy (anonymous):

Right, so all you have to do is plug it in, so let me make it look nice for you\[\frac{ (x \times y - z) }{ x } \implies \frac{ [(r+2)(2r-9)-(r^2+17r+30)] }{ r+2 }\]

OpenStudy (anonymous):

Seems reasonable, let me know if you have any doubts, can you simplify now?

OpenStudy (thedj4jc):

I'm not sure, that's how it'd be written I know. But I'm having trouble understanding how to simplify it.

OpenStudy (thedj4jc):

@iambatman